# Low-Resource Vision Challenges for Foundation Models

### 10 Jan 2024 v2

@@ 5.4 Discussion
Qualitative Results. We present qualitative results in Figure 6. Our model successfully handles cases where a certain portion of the image is a clear indication of the groundtruth. For example, dice circuits usually contain a digital number display. For historic map retrieval, correct examples have a unique coastline or river path, while in mechanical drawings the correct component is clear from all views in the drawing. However, our baselines suffer when the relationships between multiple image regions are key. For instance, the horn in the bell circuit diagram also appears in audio amplifier circuits. The mechanical drawing failure cases appear correct from one drawing perspective but not the others. Our baselines also struggle when image style is rare or unseen in training, as in the case of the Innsbruck map.
@@ 6 Conclusion
This paper studies low-resource vision tasks. We collect a benchmark of truly low-resource vision tasks and find these tasks share three common challenges: extremely limited data, fine-grained differences between images, and highly specialized domains. To combat these challenges we investigate the generalization capability of foundation models, but find they struggle on low-resource vision tasks. We thus propose three baselines, one per challenge, to make a step toward solving low-resource vision. These baselines improve over prior works which tackle individual challenges in isolation and can be easily plugged into different foundation models. Nevertheless, this is still an under-explored area with many opportunities for future work.

### 11 Apr 2024 v3

@@ 5.4 Discussion
Qualitative Results. We present qualitative results in Figure 6. Our model successfully handles cases where a portion of the image is a clear indication of the label. For example, dice circuits contain a digital number display. For historic map retrieval, correct examples have a unique coastline or river path, while in mechanical drawings the correct component is clear from all views in the drawing. However, our baselines suffer when the relationships between multiple image regions are key. For instance, the horn in the bell circuit diagram also appears in audio amplifiers. The mechanical drawing failure cases appear correct from one drawing perspective but not the others. Our baselines also struggle when the image style is rare in training, as in the Innsbruck map.
@@ 6 Conclusion
This paper studies low-resource vision. We collect a benchmark of truly low-resource vision tasks and find these tasks share three challenges: extremely limited data, fine-grained differences between images, and highly specialized domains. To combat these challenges we investigate the generalization capability of foundation models, but find they struggle on low-resource vision tasks. We thus propose three baselines, one per challenge, in a step to solving low-resource vision. These baselines improve over prior works tackling individual challenges and can be easily plugged into different foundation models. Nevertheless, low-resource vision is still under-explored with many opportunities for future work.
@@
Figure 10: Diffusion Model Thresholds have limited influence on baseline I for synthesizing data.
@@ C.1 Generated Data for Data Scarcity
Diffusion Model Threshold. Our generated data baseline is insensitive to the selection of $\gamma$ and $\tau$ thresholds and improves over ImageBind zero-shot transfer for all thresholds tested in Figure 10.
@@ C.3 Attention for Specialized Domains
Figure 15: Confusion Matrix reveals our baselines achieve stronger performance on classes with prominent patterns.
@@ E Visualizations
Confusion Matrix. We provide a confusion matrix in Figure 15. Indeed, our baselines recognize components with prominent patterns more clearly, achieving stronger performance on dice, infrared sensor, and relay. It confuses classes with shared components, e.g., audio mixer and wave generator.